The 1.49 kg head of a hammer has a speed of 6.6 m/s just before it strikes a nail and is brought to rest -Part A Estimate the temperature rise of a 12 g iron nail generated by 5.0 such hammer blows done in quick succession. Assume the nail absorbs all the energy Express your answer using two significant figures. View Available Hint(s)
Answer #1

answer) first we have to find the kinetic energy lost by hammer

so K.E=n*1/2m12v2=5*1/2*1.49kg*(6.6m/s)2=162.261J

this energy is =Q

we know the formula

Q=m2*c*dT

c=450j/kg

dT=Q/m2c=162.261J/0.012*450=30oC

so the answer is 30oC or 30.05oC